# On computational complexity and average-case hardness of shallow-depth boson sampling

### v1

@@ Appendix C Proof of Theorem 4
For the reduction from the worst-case circuit $C_{\bm{P}}$ to the average-case circuits, we sample random circuit $U(\theta)$ in $(\mathcal{B}\mathcal{B}^{*})^{q}$ with $q=q_{0}+1$, by sampling independently distributed local Haar random gate $\{H_{i}\}_{i=1}^{m}$ for gate number $m=qM\log M$, perturbing them by the Cayley transform parameterized by $\theta$ and multiplying the worst-case circuit gates from $C_{\bm{P}}$ as in Definition 7. Then, $U(\theta)$ follows the distribution $\mathcal{H}_{\mathcal{A},\theta}^{C_{\bm{P}}}$, with $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q}$, and $U(1)=C_{\bm{P}}$.
@@
Given randomly chosen outcome $\bm{s}_{\bm{P}}\sim\mathcal{G}_{M,N}$ and circuit $U(\theta)\sim\mathcal{H}_{\mathcal{A},\theta}^{C_{\bm{P}}}$, we input them in the oracle $\mathcal{O}$. For at least $1-\eta$ over $\bm{s}_{\bm{P}}$, the failure probability of $\mathcal{O}$ is at most
$\Pr_{U(\theta)\sim\mathcal{H}_{\mathcal{A},\theta}^{C_{\bm{P}}}}[|\mathcal{O}(\bm{s}_{\bm{P}},U(\theta))-p_{\bm{s}_{\bm{P}}}(U(\theta))|>\epsilon]<\delta+D_{\rm{TV}}(\mathcal{H}_{\mathcal{A},\theta}^{C_{\bm{P}}},\mathcal{H}_{\mathcal{A}}),$ (23)
@@
where $D_{\rm{TV}}$ denotes total variation distance. This is evident as we can interpret the total variation distance as the supremum over events of the difference in probabilities of those events (Viz., circuits corresponding to the failure) [32]. By Lemma 2, $D_{\rm{TV}}(\mathcal{H}_{\mathcal{A},\theta}^{C_{\bm{P}}},\mathcal{H}_{\mathcal{A}})$ is $O(m\theta)$. By setting $0\leq\theta\leq\Delta$ with $\Delta=O(m^{-1})$, we can upper bound $D_{\rm{TV}}(\mathcal{H}_{\mathcal{A},\theta}^{C_{\bm{P}}},\mathcal{H}_{\mathcal{A}})$ by an arbitrarily small constant.
@@
By Lemma 3, $p_{\bm{s}_{\bm{P}}}(U(\theta))$ is a $(4mN,4mN)$ degree rational function $\frac{P(\theta)}{Q(\theta)}$, where the denominator is given as $Q(\theta)=\left[\prod_{i=1}^{m}|q_{i}(\theta)|^{2}\right]^{N}$. We note that $Q(\theta)$ can be computed in $\Theta(m)$ time, as it only depends on the constant number of eigenvalues of local gate matrices (i.e., $\phi_{j}$ values in Eq. (4) for each local gate matrix $H_{i}$). Also, given that $\theta\leq\Delta=O(m^{-1})$, $Q(\theta)$ is very close to the unity, as $Q(\theta)\geq 1$ and
@@
Therefore, $p_{\bm{s}_{\bm{P}}}(U(\theta))$ is very close to the degree $d=4mN$ polynomial $P(\theta)$ in $\theta\in[0,\Delta]$, which allows us to use polynomial interpolation technique for $P(\theta)$. Specifically, we obtain estimations of $P(\theta)$ for different values of $\theta\in[0,\Delta]$ by querying the oracle $\mathcal{O}$, use polynomial interpolation for given $P(\theta)$ values to estimate $P(1)$, and infer the value $p_{\bm{s}_{\bm{P}}}(U(1))=p_{\bm{s}_{\bm{P}}}(C_{\bm{P}})$ by multiplying $Q(1)^{-1}$. However, $Q(1)^{-1}$ becomes arbitrarily large for the case that even a single $\phi_{i,j}$ in Eq. (C) is near ±$\pm\pi$, which will arbitrarily enlarge the imprecision of the approximation of $p_{\bm{s}_{\bm{P}}}(U(1))$. To avoid this issue, we employ the strategy from Ref. [21], which only considers the case that all $\phi_{i,j}$ values of randomly chosen gates $\{H_{i}\}_{i=1}^{m}$ are in $[-\pi+\zeta,\pi-\zeta]$, and regards the other case as failure. This happens with probability at least $1-O(m\zeta)$ over the random circuit instances. By setting $\zeta=O(m^{-1})$, we can make $O(m\zeta)$ arbitrarily small constant, and as a result, we can upper bound $Q(1)^{-1}$ (see Eq. (28) below) with high probability over random circuit instances.
Now the problem reduces to approximating degree $d=4mN$ polynomial $P(\theta)$ with the value $\mathcal{O}(\bm{s}_{\bm{P}},U(\theta))Q(\theta)$ in $\theta\in[0,\Delta]$ within additive error smaller than $\epsilon(1+O(Nm^{-1}))\approx\epsilon$; such approximations will later be used for the estimation of the value $P(1)$ via polynomial interpolation technique. The failure of $\mathcal{O}$ depends on the outcome $\bm{s}_{\bm{P}}\sim\mathcal{G}_{M,N}$ whose failure probability is at most $\eta$, and the circuit $U(\theta)\sim\mathcal{H}_{\mathcal{A},\theta}^{C_{\bm{P}}}$ whose failure probability is at most $\delta+O(m\Delta)$ from Eq. (23). Also, the probability that at least one $\phi_{i,j}$ of randomly chosen gates $\{H_{i}\}_{i=1}^{m}$ is outside of the regime $[-\pi+\zeta,\pi-\zeta]$ is at most $O(m\zeta)$. Putting everything together and applying a simple union bound, the total failure probability of the approximation of $P(\theta)$ is at most
$\begin{split}\Pr[|\mathcal{O}(\bm{s}_{\bm{P}},U(\theta))Q(\theta)-P(\theta)|>\epsilon]&<\eta+\delta+O(m\Delta)+O(m\zeta)\\ &\leq\delta^{\prime},\end{split}$ (25)
@@
Let $\{\theta_{i}\}_{i=1}^{O(d^{2})}$ be the set of equally spaced points in the interval $[0,\Delta]$. For each $\theta_{i}$, we obtain the unitary matrix $U(\theta_{i})$ using the same random gate $\{H_{i}\}_{i=1}^{m}$ and worst-case circuit $C_{\bm{P}}$. Let $y_{i}=\mathcal{O}(\bm{s}_{\bm{P}},U(\theta_{i}))Q(\theta_{i})$. By Eq. (25), each set of points $(\theta_{i},y_{i})$ satisfies
@@
From the estimated value $P(1)$, we can infer the worst-case output probability value $p_{\bm{s}_{\bm{P}}}(U(1))=P(1)/Q(1)$. As the value of $Q(1)$ depends on the values $\phi_{i,j}$ in Eq. (C), the $\phi_{i,j}$ independent lower bound of $Q(1)$ is required to set an upper bound of the additive imprecision of $p_{\bm{s}_{\bm{P}}}(U(1))$. Since we only consider the case that all of $\phi_{i,j}$ values are in $[-\pi+\zeta,\pi-\zeta]$ with $\zeta=O(m^{-1})$ for all randomly chosen gates $\{H_{i}\}_{i=1}^{m}$, we have
@@
Therefore, the total additive error for estimating $p_{\bm{s}_{\bm{P}}}(U(1))$ is bounded by $\epsilon^{\prime}2^{-2mN\log O(m^{-2})}=\epsilon 2^{O(N^{\gamma+1}(\log N)^{2})}$. By setting $\epsilon=2^{-O(N^{\gamma+1}(\log N)^{2})}2^{-O(N)}=2^{-O(N^{\gamma+1}(\log N)^{2})}$, we can estimate the worst-case output probability value $p_{\bm{s}_{\bm{P}}}(U(1))=p_{\bm{s}_{\bm{P}}}(C_{\bm{P}})=p_{\bm{s}_{0}}(C_{0})$ within additive error $2^{-O(N)}$, and the whole reduction process is in $\text{BPP}^{\text{NP}}$. This completes the proof.
For the polynomial interpolation, we employ the Robust Berlekamp-Welch algorithm recently proposed in Ref. [21].
@@
Appendix D Proof of Lemma 4

### v2

@@ Appendix D About Cayley transform
Let $\mathcal{A}$ be the circuit architecture with $m$ number of gates. For the given circuit $C$ in $\mathcal{A}$ with gates $\{G_{i}\}_{i=1}^{m}$, the circuit $V(\theta)$ is defined with each gate of $C$ replaced by $G_{i}\rightarrow H_{i}(\theta)G_{i}$, where each $H_{i}(\theta)$ is a Cayley transform of independently distributed local Haar random gate $H_{i}$ $(i\in[m])$ parameterized by $\theta\in[0,1]$ in Definition 8. We define $\mathcal{H}_{\mathcal{A},\theta}^{C}$ as the distribution for such $V(\theta)$. Here, the distribution of the $V(0)$ is the local random circuit ensemble $\mathcal{H}_{\mathcal{A}}$ in Definition 4, and $V(1)=C$.
We should make sure that the success probability of average-case approximation over circuits is still large enough after the perturbation described in Definition 9, to establish the reduction process successfully. This is evident in the case that the total variation distance over circuits induced by the perturbation is small enough, as the success probability over circuits perturbs by, at most, the total variation distance. In fact, Ref. [41] proved that total variation distance between $\mathcal{H}_{\mathcal{A}}$ and $\mathcal{H}_{\mathcal{A},\theta}^{C}$ is small for sufficiently small perturbation $\theta$ compared to the inverse gate number of the circuit architecture $\mathcal{A}$.
Lemma 8 (Movassagh [41]).
Let $\mathcal{A}$ be the circuit architecture with $m$ number of gates. For $\theta\ll 1$ and for any circuit $C$ in $\mathcal{A}$, total variation distance between $\mathcal{H}_{\mathcal{A},\theta}^{C}$ and $\mathcal{H}_{\mathcal{A}}$ is $O(m\theta)$.
Therefore, by using small $\theta=O(m^{-1})$, one can upper-bound the total variation distance by an arbitrarily small constant, which implies that the success probability of average-case approximation over circuits (i.e., $1-\delta$ in Theorem 4) also perturbs by at most a small constant.
As we have sketched in Sec. IV, we infer the worst-case output probability $p_{\bm{s}}(C\bm{P})$ for some fixed circuit $C$ and permutation matrix $\bm{P}$ via average-case output probabilities $p_{\bm{s}}(V(\theta)\bm{P})$ for perturbed random circuit $V(\theta)$. To investigate the viability of the inference of the worst-case value, we examine the behavior of the function $p_{\bm{s}}(V(\theta)\bm{P})$ characterized by the parameter $\theta$. Using Definition 8, we find that $p_{\bm{s}}(V(\theta)\bm{P})$ can be represented as a rational function in $\theta$.
Lemma 9.
Let $\mathcal{A}$ be the $q$-Kaleidoscope circuit architecture $(\mathcal{B}\mathcal{B}^{*})^{q}$ with $m=qM\log M$ number of gates, and $V(\theta)\sim\mathcal{H}_{\mathcal{A},\theta}^{C}$ for any $C$ in $\mathcal{A}$. Then for any collision-free outcome $\bm{s}$ and $M$ by $M$ permutation $\bm{P}$, the output probability $p_{\bm{s}}(V(\theta)\bm{P})$ can be represented as a degree $(4mN,4mN)$ rational funtion in $\theta$.
Proof.
For given circuit unitary matrix $V(\theta)\sim\mathcal{H}_{\mathcal{A},\theta}^{C}$ with $C$ composed of $\{G_{i}\}_{i=1}^{m}$ gates, one can decompose $V(\theta)$ with $m=qM\log M$ product of unitary matrices, such that each matrix element of $V(\theta)$ can be represented as
$\left[V(\theta)\right]_{j,k}=\sum_{l_{1}=1}^{M}\sum_{l_{2}=1}^{M}\cdots\sum_{l_{m-1}=1}^{M}V_{j,l_{1}}^{(1)}V_{l_{1},l_{2}}^{(2)}\cdots V_{l_{m-1},k}^{(m)},$ (28)
where each $V^{(i)}$ denotes an $M$-dimensional unitary matrix, with a single gate unitary matrix $H_{i}(\theta)G_{i}$ applied to the modes participating in the gate and identity for the rest of the modes. For example, if the $i$th gate $H_{i}(\theta)G_{i}$ is a two-mode gate between the first two modes, $V^{(i)}$ is a block diagonal matrix of $H_{i}(\theta)G_{i}$ and identity matrix, namely, $V^{(i)}=H_{i}(\theta)G_{i}\bigoplus I_{M-2}$.
For circuit architecture $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q}$ which is composed of only two-mode gates, matrix elements of $V^{(i)}$ can be represented as degree $(2,2)$ rational functions in $\theta$, where the common denominator for the elements is given by $q_{i}(\theta)$, defined in Eq. (26) but with index $i$ appended for $i$th random gate $H_{i}(\theta)$. Using reduction to the common denominator for all of the $m$ gates, $\left[V(\theta)\right]_{j,k}$ can be represented as $(2m,2m)$ rational function in $\theta$ with the common denominator $\prod_{i=1}^{m}q_{i}(\theta)$; note that it does not change with the indices $j,k$.
Now, by Eq. (1), the output probability $p_{\bm{s}}(V(\theta)\bm{P})$ for collision-free outcome $\bm{s}$ has the form of
$\begin{split}p_{\bm{s}}(V(\theta)\bm{P})&=\left|\sum_{\sigma\in S_{N}}\prod_{j=1}^{N}\left[V(\theta)_{\bm{s},\bm{t}^{*}}\right]_{\sigma_{j},j}\right|^{2},\end{split}$ (29)
where we denote $\bm{t}^{*}$ as an input configuration vector permuted by $\bm{P}^{-1}$ from the fixed input configuration $\bm{t}$ in Eq. (1), and $S_{N}$ is $N$-mode permutation group. One can easily check that the common denominator for $\prod_{j=1}^{N}\left[V(\theta)_{\bm{s},\bm{t}^{*}}\right]_{\sigma_{j},j}$ is $[\prod_{i=1}^{m}q_{i}(\theta)]^{N}$, and it does not change with permutation $\sigma$. Let us define $Q(\theta)=[\prod_{i=1}^{m}|q_{i}(\theta)|^{2}]^{N}$, which is a degree $4mN$ polynomial in $\theta$. Then, $Q(\theta)$ serves as the common denominator for the output probability. Hence, the output probability can be represented as $p_{\bm{s}}(V(\theta))=\frac{P(\theta)}{Q(\theta)}$, with $P(\theta)$ also a degree $4mN$ polynomial in $\theta$. ∎
Appendix E Proof of Theorem 4 (average-case hardness)
In the following, we provide a proof of Theorem 4. As we have already sketched the proof in the main text, for the sake of completeness, we will provide technical details here. Specifically, we establish the worst-to-average-case reduction: We prove that, for $\delta,\,\eta\geq 0$ with $\delta+\eta<\frac{1}{4}$, the output probability $p_{\bm{s}_{0}}(C_{0})$ in Theorem 3 can be estimated in additive imprecision $2^{-O(N)}$ in complexity class $\text{BPP}^{\text{NP}}$, given output probability estimations $p_{\bm{s}}(U)$ within additive imprecision $\epsilon=2^{-O(N^{\gamma+1}(\log N)^{2})}$, with probability at least $1-\delta$ over $U\sim\mathcal{H}_{\mathcal{A}}\mathcal{P}$ with $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q\geq 3}$ for at least $1-\eta$ over $\bm{s}\sim\mathcal{G}_{M,N}$.
Proof of Theorem 4.
Let $\mathcal{O}$ be an oracle that solves the problem in Theorem 4, i.e., on input $\bm{s}\sim\mathcal{G}_{M,N}$ and $U\sim\mathcal{H}_{\mathcal{A}}\mathcal{P}$ with $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q\geq 3}$, the oracle outputs $p_{\bm{s}}(U)$ within additive error $\epsilon$ with $1-\delta$ portion of $U$ for $1-\eta$ portion of $\bm{s}$. In other words, for $1-\eta$ of $\bm{s}\sim\mathcal{G}_{M,N}$ and $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q\geq 3}$, we have
$\Pr_{U\sim\mathcal{H}_{\mathcal{A}}\mathcal{P}}[|\mathcal{O}(\bm{s},U)-p_{\bm{s}}(U)|>\epsilon]<\delta.$ (30)
Let $C_{0}$ be the worst-case circuit in $(\mathcal{B}\mathcal{B}^{*})^{q_{0}}$ with $q_{0}\geq 1$, and let $\bm{s}_{0}$ be the fixed collision-free output given in Theorem 3. In the following, we show that approximating $p_{\bm{s}_{0}}(C_{0})$ to within additive error $2^{-O(N)}$ (i.e., solving the problem in Theorem 3) is in $\text{BPP}^{\text{NP}^{\mathcal{O}}}$, which implies that the average-case approximation of $p_{\bm{s}}(U)$ to within $\epsilon$ is #P-hard under $\text{BPP}^{\text{NP}}$ reduction.
As we have argued in the main text, we first prepare $M$ by $M$ permutation matrices $\bm{P}_{0}$ and $\bm{P}_{1}$ each sampled uniformly over the $M!$-number of possible permutations. Using these sampled permutations, we newly define permuted outcome $\bm{s}_{p}$ and permuted circuit $C_{p}$ as
$\bm{s}_{p}\coloneqq\bm{P}_{0}\bm{s}_{0},\quad C_{p}\coloneqq\bm{P}_{0}C_{0}\bm{P}_{1}^{-1},$ (31)
where one can easily check that the permuted outcome follows random outcome ensemble $\bm{s}_{p}\sim\mathcal{G}_{M,N}$ in Definition 6. Also, the permuted circuit $C_{p}$ can be implemented in the circuit architecture $(\mathcal{B}\mathcal{B}^{*})^{q_{0}+2}$, because each of the sampled $\bm{P}_{0}$ and $\bm{P}_{1}$ can be implemented in the architecture $\mathcal{B}\mathcal{B}^{*}$ by Lemma 1.
Note that, by definition in Eq. (31), the worst-case output probability can also be represented as $p_{\bm{s}_{0}}(C_{0})=p_{\bm{s}_{p}}(\bm{P}_{0}C_{0})=p_{\bm{s}_{p}}(C_{p}\bm{P}_{1})$. Hereafter, we consider $C_{p}$ as a “revised” worst-case circuit, and conduct circuit level worst-to-average-case reduction, from the worst-case circuit $C_{p}$ to the average-case circuit in $(\mathcal{B}\mathcal{B}^{*})^{q_{0}+2}$.
For the average-case circuit, we sample the perturbed random circuit $V(\theta)$ from $\mathcal{H}_{\mathcal{A},\theta}^{C_{p}}$ defined in Definition 9, for the circuit architecture $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q}$ with $q=q_{0}+2\geq 3$ from now on, and the worst-case circuit $C_{p}$. This can be done by sampling independently distributed local Haar random gates $\{H_{i}\}_{i=1}^{m}$ for gate number $m=qM\log M$, perturbing them by the Cayley transform parameterized by $\theta$ according to the gates composing the worst-case circuit $C_{p}$. Then, by Definition 9, $V(1)=C_{p}$ such that $p_{\bm{s}_{p}}(V(1)\bm{P}_{1})=p_{\bm{s}_{0}}(C_{0})$ is the worst-case output probability we aim to estimate.
Given the above $V(\theta)$, we multiply the sampled permutation $\bm{P}_{1}$ at the right side of $V(\theta)$. Because $V(0)\bm{P}_{1}$ follows the distribution $\mathcal{H}_{\mathcal{A}}\mathcal{P}$, one can expect that the distribution of $V(\theta)\bm{P}_{1}$ is close to $\mathcal{H}_{\mathcal{A}}\mathcal{P}$ for sufficiently small $\theta$. Then, one can also expect that the success probability of the oracle $\mathcal{O}$ (i.e., the right-hand side of Eq. (30)) on input a perturbed circuit $V(\theta)\bm{P}_{1}$ instead of $V(0)\bm{P}_{1}$ would not much be deviated from $\delta$ as long as $\theta$ is small enough.
@@ Appendix E Proof of Theorem 4 (average-case hardness)
Based on this understanding, we input $\bm{s}_{p}$ and $V(\theta)\bm{P}_{1}$ in the oracle $\mathcal{O}$. Then, by Eq. (30), for at least $1-\eta$ over $\bm{s}_{p}$, the failure probability of $\mathcal{O}$ is at most
$\Pr\left[\left|\mathcal{O}(\bm{s}_{p},V(\theta)\bm{P}_{1})-p_{\bm{s}_{p}}(V(\theta)\bm{P}_{1})\right|>\epsilon\right]<\delta+D_{\rm{TV}}(\mathcal{H}_{\mathcal{A},\theta}^{C_{p}},\mathcal{H}_{\mathcal{A}}),$ (32)
@@
where $D_{\rm{TV}}$ denotes total variation distance. The inequality in Eq. (32) is evident based on the two facts: We can interpret the total variation distance as the supremum over events of the difference in probabilities of those events (Viz., circuits corresponding to the failure) [40], and the equal permutation on the circuit does not affect the total variation distance over the circuit space. In the right-hand side of Eq. (32), by Lemma 8, $D_{\rm{TV}}(\mathcal{H}_{\mathcal{A},\theta}^{C_{p}},\mathcal{H}_{\mathcal{A}})$ is $O(m\theta)$. By setting $0\leq\theta\leq\Delta$ with $\Delta=O(m^{-1})$, we can upper bound $D_{\rm{TV}}(\mathcal{H}_{\mathcal{A},\theta}^{C_{p}},\mathcal{H}_{\mathcal{A}})$ by an arbitrarily small constant.
@@
By Lemma 9, $p_{\bm{s}_{p}}(V(\theta)\bm{P}_{1})$ is a $(4mN,4mN)$ degree rational function $\frac{P(\theta)}{Q(\theta)}$, where the denominator is given as $Q(\theta)=\left[\prod_{i=1}^{m}|q_{i}(\theta)|^{2}\right]^{N}$ for $q_{i}(\theta)$ defined in Definition 8 but with subindex $i$ corresponding to $i$th random gate $H_{i}(\theta)$. We note that $Q(\theta)$ can be computed in $\Theta(m)$ time, as it only depends on the constant number of eigenvalues of local gate matrices (i.e., $\phi_{j}$ values in Eq. (26) for each local gate matrix $H_{i}$). Also, given that $\theta\leq\Delta=O(m^{-1})$, $Q(\theta)$ is very close to the unity, because $Q(\theta)\geq 1$ and
@@
Therefore, $p_{\bm{s}_{p}}(V(\theta)\bm{P}_{1})$ is very close to the degree $d=4mN$ polynomial $P(\theta)$ in $\theta\in[0,\Delta]$, which allows us to use polynomial interpolation technique for $P(\theta)$. Specifically, we obtain estimations of $P(\theta)$ for different values of $\theta\in[0,\Delta]$ by querying the oracle $\mathcal{O}$, use polynomial interpolation for given $P(\theta)$ values to estimate $P(1)$, and infer the value $p_{\bm{s}_{p}}(V(1)\bm{P}_{1})=p_{\bm{s}_{0}}(C_{0})$ by multiplying $Q(1)^{-1}$. However, $Q(1)^{-1}$ becomes arbitrarily large for the case that even a single $\phi_{i,j}$ in Eq. (E) is near ±$\pm\pi$, which will arbitrarily enlarge the imprecision of the approximation of $p_{\bm{s}_{p}}(V(1)\bm{P}_{1})$. To avoid this issue, we employ the strategy from Ref. [28], which only considers the case that all $\phi_{i,j}$ values of randomly chosen gates $\{H_{i}\}_{i=1}^{m}$ are in $[-\pi+\zeta,\pi-\zeta]$, and regards the other case as failure. This happens with probability at least $1-O(m\zeta)$ over the random circuit instances. By setting $\zeta=O(m^{-1})$, we can make $O(m\zeta)$ arbitrarily small constant, and as a result, we can upper bound $Q(1)^{-1}$ (see Eq. (37) below) with high probability over random circuit instances.
Now the problem reduces to approximating degree $d=4mN$ polynomial $P(\theta)$ with the value $\mathcal{O}(\bm{s}_{p},V(\theta)\bm{P}_{1})Q(\theta)$ in $\theta\in[0,\Delta]$ within additive error smaller than $\epsilon(1+O(Nm^{-1}))\simeq\epsilon$; such approximations will later be used for the estimation of the value $P(1)$ via polynomial interpolation technique. The failure of $\mathcal{O}$ depends on the outcome $\bm{s}_{p}\sim\mathcal{G}_{M,N}$ whose failure probability is at most $\eta$, and the circuit $V(\theta)\sim\mathcal{H}_{\mathcal{A},\theta}^{C_{p}}$ whose failure probability is at most $\delta+O(m\Delta)$ from Eq. (32). Also, the probability that at least one $\phi_{i,j}$ of randomly chosen gates $\{H_{i}\}_{i=1}^{m}$ is outside of the regime $[-\pi+\zeta,\pi-\zeta]$ is at most $O(m\zeta)$. Putting everything together and applying a simple union bound, the total failure probability of the approximation of $P(\theta)$ is at most
$\Pr\left[\left|\mathcal{O}(\bm{s}_{p},V(\theta)\bm{P}_{1})Q(\theta)-P(\theta)\right|>\epsilon\right]$ $<\eta+\delta+O(m\Delta)+O(m\zeta)\leq\delta^{\prime},$ (34)
@@
Let $\{\theta_{i}\}_{i=1}^{O(d^{2})}$ be the set of equally spaced points in the interval $[0,\Delta]$. For each $\theta_{i}$, we obtain the unitary matrix $U(\theta_{i})$ using the same random gate $\{H_{i}\}_{i=1}^{m}$ and worst-case circuit $C_{p}$. Let $y_{i}=\mathcal{O}(\bm{s}_{p},V(\theta_{i})\bm{P}_{1})Q(\theta_{i})$. By Eq. (34), each set of points $(\theta_{i},y_{i})$ satisfies
@@
From the estimated value $P(1)$, we can infer the worst-case output probability value $p_{\bm{s}_{0}}(C_{0})=p_{\bm{s}_{p}}(V(1)\bm{P}_{1})=P(1)/Q(1)$. As the value of $Q(1)$ depends on the values $\phi_{i,j}$ in Eq. (E), the $\phi_{i,j}$ independent lower bound of $Q(1)$ is required to set an upper bound of the additive imprecision of $p_{\bm{s}_{p}}(V(1)\bm{P}_{1})$. Since we only consider the case that all of $\phi_{i,j}$ values are in $[-\pi+\zeta,\pi-\zeta]$ with $\zeta=O(m^{-1})$ for all randomly chosen gates $\{H_{i}\}_{i=1}^{m}$, we have
@@
Therefore, the total additive error for estimating $p_{\bm{s}_{p}}(V(1)\bm{P}_{1})$ is bounded by $\epsilon^{\prime}2^{-2mN\log O(m^{-2})}=\epsilon 2^{O(N^{\gamma+1}(\log N)^{2})}$. By setting $\epsilon=2^{-O(N^{\gamma+1}(\log N)^{2})}2^{-O(N)}=2^{-O(N^{\gamma+1}(\log N)^{2})}$, we can estimate the worst-case output probability value $p_{\bm{s}_{p}}(V(1)\bm{P}_{1})=p_{\bm{s}_{0}}(C_{0})$ within additive error $2^{-O(N)}$, and the whole reduction process is in $\text{BPP}^{\text{NP}}$. This completes the proof.
For the polynomial interpolation, we employ the Robust Berlekamp-Welch algorithm that requires the NP oracle, recently proposed in Ref. [28].
@@
Appendix F Proof of Lemma 2